OpenStudy (anonymous):

HELP!!!!!!!!!!!!!!! Simplify 4^21 ---- 4^7 1. 4^28 2. 4^21 3. 4^7 4. 4^3

OpenStudy (anonymous):

@gaby6

OpenStudy (anonymous):

@sue101

OpenStudy (anonymous):

@sunlab2015 @surdawi @MicroBot @MATHGEOMETRY8 @Hollywood_chrissy

OpenStudy (anonymous):

Always remember that : \(\frac{a^b}{a^c} = a^{b-c} \)

OpenStudy (anonymous):

Hence, \[\large{\frac{4^{21}}{4^7} = 4^{21-7} = 4^{14}}\]

OpenStudy (anonymous):

Which operation would you use on the exponents to simplify the expression 57 × 511? Addition Subtraction Multiplication Division

OpenStudy (anonymous):

Which operation would you use on the exponents to simplify the expression 5^7 × 5^11? Addition Subtraction Multiplication Division

OpenStudy (anonymous):

Addition.........

OpenStudy (anonymous):

Simplify 15^4 ---- 15^5

OpenStudy (anonymous):

@Miyuru

OpenStudy (anonymous):

@SheldonEinstein

OpenStudy (anonymous):

Do subtraction.. :)

OpenStudy (anonymous):

Simplify 15^4 ---- 15^5

OpenStudy (anonymous):

wouldnt it be 15^-9??? not a answer

OpenStudy (anonymous):

its 15^-4 and 15^5

OpenStudy (anonymous):

Do subtraction like this....\[\frac{15^4}{15^5}=15^{4-5}=15^{-1}\]

OpenStudy (anonymous):

This answer is for the one you typed before..

OpenStudy (anonymous):

Addition

OpenStudy (radar):

\[15^{(4-5)}\]

OpenStudy (anonymous):

\[\frac{15^{-4}}{15^5}=15^{-4-5}=15^{-9}\]